Assume labor and machines are used in the production of an output.   The marginal product of the last worker hired was 175 units a day and the marginal product of the last machine employed was 850 units a day. The wage to hire a worker for a day is $100 while the cost to employ the machine for a day was $400. Assuming perfect competition what would you recommend the firm do to become more efficient?
Answer #1

Answer

reduce labor and increase machines up to MP per dollar is equal.

========

A firm maximizes profit at hiring where marginal product per dollar is equal from both inputs.

MP per dollar =MP/price of the input

MP per dollar of labor =175/100=1.75

MP per dollar of machine =850/400=2.125

the firm should reduce the labor input and increase the machines as the MP per dollar of a machine is higher than the labor.